# footnotetext: ${}^{\text{\Letter}}$ Corresponding Authors

### v1

@@ Abstract
In this paper, we investigate the thermodynamics especially the Joule-Thomson expansion of the $n$-dimensional FRW (Friedmann-Robertson-Walker) universe with a perfect fluid. We derive the thermodynamic equations of state $P=P(V,T)$ for the $n$-dimensional FRW universe in Einstein gravity and Einstein-Gauss-Bonnet gravity, where the thermodynamic pressure $P$ defined by the work density $W$ of the perfect fluid, $i.e.$ $P\equiv W$. Furthermore, we present the Joule-Thomson expansion as an application of these equations of state to elucidate the cooling-heating properties of the $n$-dimensional FRW universe. We determine the inversion temperature and inversion pressure in the FRW universe with arbitrary dimensions for the first time, and illustrate the characteristics of inversion curves and isenthalpic curves in the $T$-$P$ plane. This study offers insights into deepening our comprehension of cooling and heating regions in the FRW universe, thereby revealing its expansion mechanisms.
@@ I Introduction
Since the discovery of Hawking radiation Hawking:1974sw and the first proposal of Bekenstein-Hawking entropy Bardeen1973 ; Bekenstein:1974ax , people found that black hole can be viewed as thermodynamic system, making it possible to have some interesting thermodynamic similarities with classical thermodynamic systems. In the presence of cosmological constant, Hawking and Page Hawking:1982dh has investigated the thermodynamics of Schwarzschild-AdS (anti-de Sitter) black hole, and found an interesting phase transition between the black hole and thermal gas in AdS space. Since then, significant attention has been devoted to the study of asymptotically AdS black holes because they admit gauge gravity duality description Witten:1998zw . In the previous studies, the cosmological constant is treated as a fixed parameter. Studing the thermodynamic properties of black holes in AdS space is of great importance to the understanding of the AdS/CFT correspondence. In recent years, people treat the negative cosmological constant $\Lambda$ as the thermodynamic pressure $P=-{\Lambda}/{8\pi}$ Kastor:2009wy ; Dolan:2010ha , and found many other interesting thermodynamic phenomena in the extended phase space of AdS black holes, such as $P$-$V$ criticality Kubiznak:2012wp ; Cai:2013qga ; Xu:2015rfa ; Hu:2018qsy ; Abdusattar:2023xxs , Joule-Thomson expansion Okcu:2016tgt ; Feng:2020swq ; Barrientos:2022uit , etc.
@@
An interesting question is that whether the thermodynamic properties of the FRW universe share some similarities with black holes. A remarkable feature of thermodynamics is its universality. The largest known system controlled by gravity, namely the universe, should also obey the laws of thermodynamics. According to the spirit of Jacobson Jacobson:1995ab that the Einstein field equation can be derived from the Clausius relation, Cai and Kim Cai:2005ra discovered that Hawking radiation also exist in the FRW universe, and studied the Friedmann equation of the FRW universe at the apparent horizon. Later, the unified first law of the FRW universe at the apparent horizon was studied in Cai:2006rs , and similar connections were also found in Einstein-Gauss-Bennet gravity Akbar:2006kj ; Akbar:2006er and others.
@@
As an extension, we have recently conducted first principle study of thermodynamic pressure for the four-dimensional FRW universe thermodynamics Abdusattar:2021wfv . We investigated the thermodynamics for the FRW universe, especially presented the thermodynamic equation of state for the four dimensional FRW universe in Einstein gravity, $i.e.$ $P=P(R_{A},T)$ for the first time, where the thermodynamic pressure $P$ is defined by the work density $W$ of the perfect fluid Hayward:1993wb ; Hayward:1994bu
@@
here $h_{ij}$ and $T^{ij}$ are the $0,1$-components of the metric and the stress-tensor Akbar:2006kj with $i,j=0,1,x^{0}=t,x^{1}=r$. In addition, we extended the thermodynamic properties of the FRW universe to modified theories of gravity and found that the $P$-$V$phase transitions Kong:2021qiu ; Abdusattar:2023hlj . It is worth noting that the coexisting phase of the $P$-$V$ phase transition in the FRW universe is higher than the critical temperature, which distinguishes it from van der Waals systems and most of AdS black holes system. It enables us to study the micro-structure of the FRW universe as a thermodynamic system Abdusattar:2023pck . Our analysis of the Ruppeiner geometry for the FRW universe provided new insights into the nature of interactions among the perfect fluid matter constituents in the expanding FRW universe from the viewpoint of geometro-thermodynamics. Subsequently, encouraged by the methods in the above works, the thermodynamic properties of the FRW universe with same method has also been studied in some of modified theories of gravity Saavedra:2023lds ; Housset:2023jcm .
@@
Furthermore, as an application of thermodynamic equations of state, we study the Joule-Thomson expansion for the $n$-dimensional FRW universe. In classical thermodynamics, the Joule-Thomson expansion process of a van der Waals system describes a gas moving from high pressure region to low pressure region at a fixed enthalpy. In Okcu:2016tgt , the author creatively applied the famous Joule-Thomson expansion to the charged AdS black hole and found the existence of inversion temperature and inversion pressure. Subsequently, various black holes were studied, such as Gauss-Bonnet black holes Lan:2018nnp , regular (Bardeen)-AdS black hole Pu:2019bxf , Born-Infeld AdS black holes Bi:2020vcg , hairy black holes Xing:2021gpn , nonlinearly charged AdS black holes Kruglov:2022lnc , see for Ghaffarnejad:2020nhb ; Nam:2019zyk ; Dehghani:2023yph ; Du:2023kgj ; Feng:2022wnq more related works. In our previous work, we have studied the Joule-Thomson expansion of van der Waals system to the four-dimensional FRW universe under Einstein gravity Abdusattar:2021wfv and  [^1] In the braneworld scenario, our universe is modeled as a brane with three spatial dimensions embedded in a higher dimensional bulk, see for Meehan:2014zsa ; Abdusattar:2015azp related works and the references therein.braneworld scenario Kong:2022xny , hence found the existence of inversion temperature and inversion pressure. As a generalization, we extend the research on the Joule Thomson expansion of the FRW universe to the Einstein gravity and Einstein-Gauss-Bonnet gravity models with arbitrary dimensions. We introduce the $T$-$P$ isenthalpic expansion process in a high-dimensional FRW universe, pioneering the exploration of thermodynamic properties in arbitrary dimensions for the first time. Drawing upon these investigations, we aim to deduce the thermodynamic characteristics of the higher-dimensional FRW universe and further assess how Einstein-Gauss-Bonnet gravity modifies these properties, offering a comprehensive understanding of the thermodynamic behavior in such systems.
@@ II.1 Apparent Horizon and Hawking Temperature
where $a(t)$ is the time-dependent scale factor, $k=-1$, $+1$, 00 are the spatial curvature corresponds to an open (Hyperbolic), closed (Spherical), and flat universe, and $d\Omega_{(n-2)}^{2}$ represents the metric of the $(n-2)$-dimensional unit sphere.
For convenience later discussion, we rewrite the line-element of $n$-dimensional FRW universe (2) with physical radius $R\equiv a(t)r$ in the following form
@@
where $H\equiv\dot{a}/a$ is the Hubble parameter, and ``"``"`$\cdot$"` ` ⋅ " stands for the derivative with respect to the cosmic time. The above equation (5) shows the time dependence of the apparent horizon.
@@
Assuming $\dot{R}_{A}$ to be a small quantity, the surface gravity $\kappa$ of the apparent horizon in the FRW universe is negative [^2] See relevant previously works on negative surface gravity in Refs. Abdusattar:2021wfv ; Hayward:1993wb ; Abdusattar:2022bpg ; Dolan:2013ft .. The Hawking temperature of the FRW universe is defined from the surface gravity given (6)
@@ II.2 The Equation of State of Arbitrary Dimensional FRW Universe
where $\mathscr{R}$ is the $n$–dimensional Ricci scalar, $k_{n}\equiv\sqrt{8\pi G_{n}}$, in which $G_{n}$ is the $n$-dimensional gravitational constant, $\alpha$ is a constant with unit of length squared and $S_{m}$ is the action of the matter fields. From Eq. (8), the Einstein field equation is given by
@@
Applying the field equations (9) to the line-element (2) with the Eq.(10) one gives the Friedmann’s equations
@@
Differentiating the above equation, we easily obtain the following relation [^3] Note that the minus sign before $TdS$ in (16) is caused by the negative surface gravity, $i.e.$ $\kappa<0$, so keeping the corresponding temperature $T$ positive will impose constraints on the perfect fluid. The relevant derivation is shown in the Appendix (60).
$dE=-TdS+WdV\,,$ (16)
@@
with $A:=V_{n-2}^{1}R_{A}^{n-2}$ is the area associated with apparent horizon, where $V_{n-2}^{1}$ is the surface area of the ($n-2$)-dimensional unit Einstein space defined by $V_{n-2}^{1}:=2\pi^{(n-1)/2}/\Gamma[{(n-1)}/{2}]$, and $V$ is the thermodynamic volume
@@
we see that the internal energy $U$ and thermodynamic pressure $P$ can be identified with $-E$ and $W$, $i.e.$
@@
The thermodynamic equation of state for the $n$-dimensional FRW universe in Einstein gravity from Eqs.(21), (7) and (14) obtained by  [^4] From now on, we set $G_{n}=1$ for simplicity in this section.
@@
The stability of the thermodynamic system can be understood by investigating the specific heat capacity at constant pressure $C_{P}$, which is positive for a stable system and negative for an unstable system. By using the Eqs.(17) and (22), one can obtain the heat capacity of the $n$-dimensional FRW universe within the framework of Einstein gravity
@@
where ${\cal H}\equiv U+PV$ is the enthalpy of the system, which derivation is given in Appendix (61). Clearly, we see that the heat capacity $C_{P}$ is always positive since the temperature should take positive value, implying the universe in this condition is thermally stable. Note that the $C_{P}$ recovers to the result in our previous work, when $n=4$ Abdusattar:2021wfv .
@@ III.1 A Brief Introduction of Einstein-Gauss-Bonnet Gravity and Friedmann’s Equation
possesses topological invariance within the realm of four-dimensional spacetimes. Notably, when integrated over a closed four-dimensional manifold without boundary, it serves as a topological invariant. Furthermore, Gauss-Bonnet gravity gracefully emerges as the second-order term within Lovelock gravity, maintaining the esteemed property that its equations of motion exclusively involve second-order derivatives of the metric. This elegant feature parallels the simplicity and elegance observed in Einstein gravity (Lovelock:1971yv, ).
@@
being the contribution coming from the Gauss-Bonnet term (27), the so–called Lanczos tensor. As we know, for $n=4$, the tensor $H_{\mu\nu}$ is zero because the Gauss-Bonnet term (27) becomes a topological invariant, not affecting the dynamic of the theory.
@@ III.2 The Equation of State of Arbitrary Dimensional FRW Universe
where $V$ is the conjugate thermodynamic volume of $W$ and can still take the form of (18), $T$ and $W$ are given in (7) and (33), respectively. Note that the minus sign before $TdS$ in (35) arises from the fact that the surface gravity $\kappa$ on apparent horizon is negative [^6] See more related discussions in the literature Akbar:2006kj . in the FRW universe, $i.e.$ $\kappa<0$, which will impose a constraint on a perfect fluid in the framework of Einstein-Gauss-Bonnet gravity obtained by
$1+\frac{p}{\rho}<\frac{4}{n-1}\Big{(}1+\frac{\widetilde{\alpha}}{R_{A}^{2}+\widetilde{\alpha}}\Big{)}\,,$ (36)
see Appendix (59) for related derivation in detail.
The entropy $S$ for the $n$-dimensional FRW universe in Einstein-Gauss-Bonnet gravity is given by (Cai:2009ua, ; Mukherji:2002de, ; Kaul:2000kf, ; Fernandes:2020rpa, )
@@
In the same way as we did for Einstein gravity, here again one identifies the internal energy $U$ with $-E$, $i.e.$ $U\equiv-E$, and the thermodynamic pressure $P$ with $W$, $i.e.$
$P\equiv$ $W\,,$ (38)
then the standard form of thermodynamic first law can be established $dU=TdS-PdV$.
After determining the definition of thermodynamic pressure, using Eqs.(7), (33) and (38), we further obtain the equation of state for the FRW universe in Einstein-Gauss-Bonnet gravity, $i.e.$ [^7] If $\widetilde{\alpha}<0$, the equation of state for $R_{A}=\sqrt{-2\widetilde{\alpha}}$ is independent of temperature, refering to a thermodynamic singularity. This behaviour is the same as in the modified gravity with a generalized conformal scalar field Kong:2021qiu ; Abdusattar:2023pck . If the Gauss-Bonnet term is absent or $\widetilde{\alpha}=0$, it simplifies to the result obtained in Einstein gravity (22). In the following discussion of this section, we will set $G_{n}=1$ for simplicity.
@@
The stability of the thermodynamic system can be understood by investigating the specific heat capacity at constant pressure $C_{P}$, which is positive for a stable system and negative for a unstable system. By using the Eqs.(37) and (39), one can obtain the heat capacity of the $n$-dimensional Gauss-Bonnet FRW universe
@@
We see that the sign of the denominator determines the stability of the universe because the numerator is always positive. A positive denominator implies a stable universe with positive heat capacity $C_{P}>0$, while a negative denominator implies an unstable universe with negative $C_{P}<0$.

### v2

@@ Abstract
In this paper, we investigate the thermodynamics especially the Joule-Thomson expansion of the $n$-dimensional FRW (Friedmann-Robertson-Walker) universe with a perfect fluid. We derive the thermodynamic equations of state $P=P(V,T)$ for the $n$-dimensional FRW universe in Einstein gravity and Einstein-Gauss-Bonnet gravity, where the thermodynamic pressure $P$ is defined by the work density $W$ of the perfect fluid, $i.e.$ $P\equiv W$. Furthermore, we present the Joule-Thomson expansion as an application of these equations of state to elucidate the cooling-heating properties of the $n$-dimensional FRW universe. We determine the inversion temperature and inversion pressure in the FRW universe with arbitrary dimensions for the first time, and illustrate the characteristics of inversion curves and isenthalpic curves in the $T$-$P$ plane. We also examine constraints on the perfect fluid in the FRW universe, as derived from the cooling-heating transition point. This study offers insights into deepening our comprehension of cooling and heating regions in the FRW universe, thereby revealing its expansion mechanisms.
@@ I Introduction
Since the discovery of Hawking radiation Hawking:1974sw and the first proposal of Bekenstein-Hawking entropy Bardeen1973 ; Bekenstein:1974ax , people have found that black holes can be viewed as thermodynamic systems, exhibiting interesting thermodynamic similarities to classical thermodynamic systems. Subsequent studies generalized this entropy concept to incorporate quantum corrections, gravitational effects in non-static spacetimes, and even cosmological horizons, leading to the framework of generalized entropy Bousso:1999xy ; Wall:2010jtc – a key step linking black hole thermodynamics to broader gravitational systems. In the presence of a cosmological constant, Hawking and Page Hawking:1982dh investigated the thermodynamics of Schwarzschild-AdS (anti-de Sitter) black holes and found an interesting phase transition between the black hole and thermal gas in AdS space. Since then, significant attention has been devoted to the study of asymptotically AdS black holes because they admit a gauge-gravity duality description Witten:1998zw . This duality has not only deepened the understanding of black hole thermodynamics but also enabled its extension to holographic cosmology, where the entropy and thermodynamics of cosmological spacetimes ($e.g.,$ FRW universe) are holographically encoded in boundary quantum field theories Maldacena:1997re ; Banks:2001px ; Cai:2005ra . In previous studies, the cosmological constant was treated as a fixed parameter. Studying the thermodynamic properties of black holes in AdS space is of great importance for understanding the AdS/CFT correspondence. In recent years, researchers have treated the negative cosmological constant $\Lambda$ as the thermodynamic pressure $P=-\Lambda/(8\pi)$ Kastor:2009wy ; Dolan:2010ha and discovered many other interesting thermodynamic phenomena in the extended phase space of AdS black holes, such as $P$-$V$ criticality Kubiznak:2012wp ; Cai:2013qga ; Xu:2015rfa ; Hu:2018qsy ; Abdusattar:2023xxs ; Abdusattar:2024zzi , Joule-Thomson expansion Okcu:2016tgt ; Feng:2020swq ; Barrientos:2022uit , etc.
@@
An interesting question is that whether the thermodynamic properties of the FRW universe share some similarities with black holes. A remarkable feature of thermodynamics is its universality. The largest known system controlled by gravity, namely the universe, should also obey the laws of thermodynamics. According to the spirit of Jacobson Jacobson:1995ab that the Einstein field equation can be derived from the Clausius relation, Cai and Kim Cai:2005ra discovered that Hawking radiation also exist in the FRW universe, and studied the Friedmann equation of the FRW universe. Later, the unified first law of the FRW universe was also studied Cai:2006rs , which after being projected onto the apparent horizon leads to the first law of thermodynamics. It has been shown that the Clausius relation holds for the FRW universe in Einstein gravity and some modified theories of gravity Akbar:2006kj ; Akbar:2006er - a result that also suggests the FRW universe is an (quasi-)equilibrium thermodynamic system Cai:2006pa ; Wu:2009wp .
@@
As an extension, we have recently conducted first principle study of thermodynamic pressure for the four-dimensional FRW universe thermodynamics Abdusattar:2021wfv . We investigated the thermodynamics for the FRW universe, especially presented the thermodynamic equation of state for the four dimensional FRW universe in Einstein gravity, $i.e.$ $P=P(V,T)$ for the first time, where the thermodynamic pressure $P$ is defined by the work density $W$ of the perfect fluid Hayward:1993wb ; Hayward:1994bu
@@
here $h_{ij}$ and $T^{ij}$ are the $0,1$-components of the metric and the stress-tensor Akbar:2006kj with $i,j=0,1,x^{0}=t,x^{1}=r$. In addition, we extended the thermodynamic properties of the FRW universe to modified theories of gravity and found that the $P$-VVphase transitions Kong:2021qiu ; Abdusattar:2023hlj . It is worth noting that the coexisting phase of the $P$-$V$ phase transition in the FRW universe is higher than the critical temperature, which distinguishes it from van der Waals systems and AdS black holes system. It enables us to study the micro-structure of the FRW universe as a thermodynamic system Abdusattar:2023pck . Our analysis of the Ruppeiner geometry for the FRW universe provided new insights into the nature of interactions among the perfect fluid matter constituents in the expanding FRW universe from the viewpoint of geometro-thermodynamics. Subsequently, encouraged by the methods in the above works, the thermodynamic properties of the FRW universe with same method has also been studied in some of modified theories of gravity Saavedra:2023lds ; Housset:2023jcm .
@@
Furthermore, as an application of thermodynamic equations of state, we study the Joule-Thomson expansion for the $n$-dimensional FRW universe. In classical thermodynamics, the Joule-Thomson expansion process of a van der Waals system describes a gas moving from high pressure region to low pressure region at a fixed enthalpy. In Okcu:2016tgt , the author creatively applied the famous Joule-Thomson expansion to the charged AdS black hole and found the existence of inversion temperature and inversion pressure. Subsequently, various black holes were studied, such as Gauss-Bonnet black holes Lan:2018nnp , regular (Bardeen)-AdS black hole Pu:2019bxf , Born-Infeld AdS black holes Bi:2020vcg , hairy black holes Xing:2021gpn , nonlinearly charged AdS black holes Kruglov:2022lnc , see Ghaffarnejad:2020nhb ; Nam:2019zyk ; Dehghani:2023yph ; Du:2023kgj ; Feng:2022wnq for more related works. In our previous work, we have studied the Joule-Thomson expansion of the four-dimensional FRW universe under Einstein gravity Abdusattar:2021wfv and braneworld scenario Kong:2022xny  [^1] In the braneworld scenario, our universe is modeled as a brane with three spatial dimensions embedded in a higher dimensional bulk, see for Meehan:2014zsa ; Abdusattar:2015azp related works and the references therein., hence found the existence of inversion temperature and inversion pressure. As a generalization, we extend the research on the Joule-Thomson expansion of the FRW universe to the Einstein gravity and Einstein-Gauss-Bonnet gravity models with arbitrary dimensions. We introduce the $T$-$P$ isenthalpic expansion process in a high-dimensional FRW universe, pioneering the exploration of thermodynamic properties in arbitrary dimensions for the first time. Drawing upon these investigations, we aim to deduce the thermodynamic characteristics of the higher-dimensional FRW universe and further assess how Einstein-Gauss-Bonnet gravity modifies these properties, offering a comprehensive understanding of the thermodynamic behavior in such systems.
@@ II.1 Apparent Horizon and Hawking Temperature
where $a(t)$ is the time-dependent scale factor, k=1k=-1, +1+1, 0 is the spatial curvature corresponding to an open (hyperbolic), closed (spherical), and flat universe, and $d\Omega_{(n-2)}^{2}$ represents the metric of the (n2)(n-2)-dimensional unit sphere.
For convenience of later discussion, we rewrite the line-element of $n$-dimensional FRW universe (2) with physical radius $R\equiv a(t)r$ in the following form
@@
where $H\equiv\dot{a}/a$ is the Hubble parameter, and $\cdot$" stands for the derivative with respect to the cosmic time. The above equation (5) shows the time dependence of the apparent horizon.
@@
Assuming $\dot{R}_{A}$ to be a small quantity, the surface gravity $\kappa$ of the apparent horizon in the FRW universe is negative [^2] See relevant previously works on negative surface gravity in Refs. Abdusattar:2021wfv ; Hayward:1993wb ; Abdusattar:2022bpg ; Dolan:2013ft .. The Hawking temperature of the FRW universe is defined from the surface gravity (6)
@@
Here, we assume that $\dot{R}_{A}\ll 1$ and concentrate on the situation of approximate thermal equilibrium, where the FRW universe undergoes an extremely gradual evolution, resulting in all thermodynamic processes being quasistatic.
@@ II.2 The Equation of State of Arbitrary Dimensional FRW Universe
where $\mathscr{R}$ is the nn–dimensional Ricci scalar, $k_{n}\equiv\sqrt{8\pi G_{n}}$, in which $G_{n}$ is the $n$-dimensional gravitational constant, $S_{m}$ is the action of the matter fields. From Eq. (8), the Einstein field equation is given by
@@
Applying the field equations (9) to the line-element (2) with the Eq.(10) one gets the Friedmann’s equations
@@
Differentiating the above equation, we easily obtain the following relation
dE=TdS+WdV,dE=-TdS+WdV\,, (16)
@@
with $A:=V_{n-2}^{1}R_{A}^{n-2}$ is the area associated with apparent horizon, where $V_{n-2}^{1}$ is the surface area of the (n2n-2)-dimensional unit Einstein space defined by $V_{n-2}^{1}:=2\pi^{(n-1)/2}/\Gamma[{(n-1)}/{2}]$, and $V$ is the thermodynamic volume
@@
we see that the internal energy $U$ and thermodynamic pressure $P$ can be identified with E-E and $W$, $i.e.$
@@
The thermodynamic equation of state for the $n$-dimensional FRW universe in Einstein gravity from Eqs.(21), (7) and (14) is obtained by [^3] From now on, we set $G_{n}=1$ for simplicity in this section.
@@
Thermodynamic stability means a system’s ability to stay in or return to a balanced state when small changes occur. For instance, the heat capacity at constant pressure $C_{P}$ is a key indicator: a positive $C_{P}$ corresponds to a stable system, while a negative $C_{P}$ indicates an unstable one. By using Eqs. (17) and (22), we can derive the heat capacity of the $n$-dimensional FRW universe within the framework of Einstein gravity
@@
where ${\cal H}\equiv U+PV$ is the enthalpy of the system, and its specific expression is given in Appendix (69). Clearly, the heat capacity $C_{P}$ is always positive, which indicates that the universe governed by Einstein gravity is thermodynamically stable. It is worth noting that when $n=4$, the value of $C_{P}$ is consistent with the result obtained in our previous work Abdusattar:2021wfv , and this consistency also reflects the stability. This thus demonstrates that the thermodynamic stability of the FRW universe under Einstein gravity holds steady, irrespective of variations in spacetime dimension.
@@ III.1 A Brief Introduction of Einstein-Gauss-Bonnet Gravity and Friedmann’s Equation
possesses topological invariance within the realm of four-dimensional spacetimes. Notably, Gauss-Bonnet gravity gracefully emerges as the second-order term within Lovelock gravity, maintaining the esteemed property that its equations of motion exclusively involve second-order derivatives of the metric. This elegant feature parallels the simplicity and elegance observed in Einstein gravity (Lovelock:1971yv, ).
@@
being the contribution coming from the Gauss-Bonnet term (27). As we know, for $n=4$, the tensor $H_{\mu\nu}$ is zero because the Gauss-Bonnet term (27) transforms into a topological invariant, which does not affect the dynamics of the theory.
@@ III.2 The Equation of State of Arbitrary Dimensional FRW Universe
where $V$ is the conjugate thermodynamic volume of $W$ and can still take the form of (18), $T$ and $W$ are given in (7) and (33), respectively. Note that the minus sign before $TdS$ in (35) arises from the fact that the surface gravity $\kappa$ on apparent horizon is negative in the FRW universe, $i.e.$ $\kappa<0$ Akbar:2006kj , which will impose a constraint on a perfect fluid in the framework of Einstein-Gauss-Bonnet gravity, see Appendix A for related discussions.
The entropy $S$ for the $n$-dimensional FRW universe in Einstein-Gauss-Bonnet gravity is given by (Cai:2005ra, ; Akbar:2006kj, ) [^5] Entropy of stationary black holes can be derived using methods such as the Euclidean or Noether charge approach Iyer:1995kg . Wald proposed a local geometric approach for dynamical black holes Wald:1993nt , which Hayward further adapted using the Kodama vector Hayward:1998ee . Although this approach could potentially be adapted to the FRW universe within the context of this modified gravity, it is notably more intricate and will therefore not be pursued in this paper.
@@
In the same way as we did for Einstein gravity, here again we compare the Eq.(35) with the standard form of thermodynamic first law Hayward:1997jp ; Wu:2009wp
$dU=TdS-PdV\,$, (37)
we see that the internal energy $U$ is identified with E-E, $i.e.$ $U\equiv-E$, and the thermodynamic pressure $P$ with $W$, $i.e.$
$P\equiv$ $W\,.$ (38)
After determining the definition of thermodynamic pressure, using Eqs.(7), (33) and (38), we further obtain the equation of state for the FRW universe in Einstein-Gauss-Bonnet gravity, i.e.i.e. [^6] If $\widetilde{\alpha}<0$, the equation of state for $R_{A}=\sqrt{-2\widetilde{\alpha}}$ is independent of temperature, refering to a thermodynamic singularity. This behaviour is the same as in the modified gravity with a generalized conformal scalar field Kong:2021qiu ; Abdusattar:2023pck ; Abdusattar:2024alq . If the Gauss-Bonnet term is absent or $\widetilde{\alpha}=0$, it simplifies to the result obtained in Einstein gravity (22). In the following discussion of this section, we will set $G_{n}=1$ for simplicity.
@@
As in the previous section, using Eqs. (36) and (39), we can calculate the heat capacity of the $n$-dimensional Gauss-Bonnet FRW universe
@@
We see that the sign of the denominator determines the stability of the universe because the numerator is always positive. A positive denominator implies a stable universe with positive heat capacity $C_{P}>0$, while a negative denominator implies an unstable universe with negative $C_{P}<0$. The expression of the heat capacity (40) can be also used to conveniently calculate the Joule-Thomson coefficient as shown in the following section.